OpenStudy (anonymous):

(5+√(2))(5-√2) multiply

OpenStudy (anonymous):

(5+√(2))(5-√2) =25-5√2+5√2-2 =23

OpenStudy (amistre64):

conjugates are the factored form of a difference of squares

OpenStudy (amistre64):

(a^2 - b^2) <=> (a-b)(a+b)